a pipe open at both ends produce a note of frequency f1 when pipe is kept with 3/4 of its lenght in water it produces a note of frequency is f2 what is ratio b/w f1 and f2a) 1/2b) 1/4c)3/4d)Not

Grade:12th pass

1 Answers

Phanindra
6 years ago
When the pipe is in air it acts as open pipe so f1 igiven by f1=v/2l...here l =resonating length.....when it is immersed in waterof 3l/4 then resonating length is l/4...so f2 = v/4×l/4...so f2 = v/l
So ratio is f1/f2 = 1/2 ...(a)...